MOA-II Obsevation Log
Return to index
Previous day
Next day
DATE: 2009/06/03
OBSERVER: Miyake Suzuki Paul
CONDITION: clear
COMMENT: took V-band image for LMC002(lmc2), MB09273 have anomaly, We could not get GOOD DARK(5360, 5364)
Compressor gas pressure: 1.5 MPa
Expected Observation Time: START 18:21 END 07:11
Go to the Seeing Log
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 18847-18886 | 40 | 372 | 22320 |
LMC | 4070-4072 | 3 | 6 | 1800 |
SMC | 837-837 | 1 | 2 | 600 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 44 | 380 | 24720 (06h52m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
13 | 13 | 38 | 38 | 39 | 1 | 6 | 13 | 39 | 38 | 7 | 6 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
15 | 37 | 6 | 6 | 24 | 13 | 7 | 6 | 6 | 1 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
0 | 3 | 0 | 0 | 0 | 0 | 1 | 1 | 1 | 0 | 0 | 0 |
Field13 | Field14 | Field15 | Field16 |
0 | 0 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
1 | 1 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAMEBR>
CHECK SHEET
Time | Status | Weather |
18:21 | observing | clear |
19:00 | observing | clear |
20:00 | observing | clear |
21:00 | observing | clear |
22:00 | observing | clear |
23:00 | observing | clear |
00:00 | observing | clear |
01:00 | observing | clear |
02:00 | observing | clear |
03:00 | observing | clear |
04:00 | observing | clear |
05:00 | observing | clear |
06:00 | observing | clear |
07:00 | observing | clear |
07:11 | observing | clear |
CCD TEMPERATURE
Previous day
Next day

VACUUM
Previous day
Next day

TELESCOPE TEMPERATURE
Previous day
Next day

Return to index
Previous day
Next day